# Conformal four-point ladder integrals in diverse dimensions and polylogarithms

### v1

@@ 4. The case $D=10$ ($\lambda=4$) and indices on the lines $\alpha=\beta=1$, $\gamma=4$:
Two-dimensional factorization for generic $\beta$
In the subsection 2.2 we concluded that the $D$-dimensional conformal ladder diagram with the generic index $\beta$ for $D>2$ and $D=2$ are respectively presented in the form (2.32) and (2.33). We also explicitly showed that the representation (2.32) admits a dimensional shift with the help of the operator (2.35) [76, 73, 74, 75], namely
@@
An amazing consequences of these results is that we can, following the logic used in works [76, 74, 73, 75] (and in subsection 2.3 to construct operator (2.47)), reduce the consideration for the arbitrary even dimension case to the two-dimensional case ($\lambda=0$). This can be seen as a huge advantage, since the two-dimensional integrals are known to have very constraint form (see e.g. [90]). In particular, we focus on the results of the work [27], where the notable factorization in $z$ and $\bar{z}$ was shown [^6] Note the similar factorization for the two-dimensional two-loop master diagram [19]..
@@
In the case $\lambda=0$ and generic $\beta$, the following representations for the function $\widetilde{\Phi}^{(\beta)}_{L}(z,\bar{z};0)$ holds [^7] In the work [76] representation (3.31) is called Fourier-Mellin representation.
@@
Proof. The first equality in (3.31) follows from the representation (2.34). To prove (3.32) we use the result from Appendix D
@@
which leads to the (3.32). ∎
Remark 10. Combining the result of proposition 1 (see eq. (2.36)) and representation (3.32) we obtain the result for arbitrary $\lambda\in\mathbb{Z}_{>0}$ and generic $\beta$ based on two-dimensional factorized formula
@@
Remark 11. In the case $\beta\in\mathbb{Z}$ the factor contained sines can be simplified to exclude the dependence on $\varepsilon$, namely
@@
Consideration of conformal ladder diagrams for $\beta=2,3,\ldots$
@@
In this subsection we consider the case of special discrete points $\beta=2,3,\ldots$. The expression $\Phi^{(\beta)}_{L}({\sf x},{\sf y})$ (2.16) for the ladder diagram is a function of the complex variable $\beta$ and this function is defined, in a standard way, as an analytical continuation from the region $0<\operatorname{Re}\beta<D/2$ where all integrals in $\Phi^{(\beta)}_{L}({\sf x},{\sf y})$ converge absolutely. This region is determined by two conditions. The horizontal line on Fig. 2 has index $D/2-\beta$ and the corresponding singularity of the integrand is integrable under the condition $D/2-\operatorname{Re}\beta<D/2$, which results in $\operatorname{Re}\beta>0$. The vertical line has index $\beta$ and the singularity is integrable under the condition $\operatorname{Re}\beta<D/2$. Combining these conditions we obtain the region of convergence $0<\operatorname{Re}\beta<D/2$. In this region of complex variable $\beta$ the function $\Phi^{(\beta)}_{L}({\sf x},{\sf y})$ is regular. The analytical continuation from this region is meromorphic function which has poles at the points $\beta=D/2+k=\lambda+1+k$, where $k=0,1,2,\ldots$.
@@
Let us start from the case $D=2$ and use (2.33) to demonstrate the origin of singularities at discrete set of points $\beta=1,2,3,\ldots$. The contour of integration (see Remark 2) separates two series of poles of the integrand: the first sequence at the points $\nu_{k}=-{\sf i}\frac{n+1}{2}-{\sf i}k,k=0,1,\ldots$ and the second sequence at the points $\nu_{k}=-{\sf i}\beta+{\sf i}\frac{n+1}{2}+{\sf i}k,k=0,1,\ldots$. For real $\beta$ and $0<\beta<1$ such a contour exists for all $n\in\mathbb{Z}_{\geq 0}$. But for $\beta\to 1$ one obtains the pinch of the contour by two poles at $\nu=-{\sf i}\frac{1}{2}$ and at $\nu={\sf i}\frac{1}{2}-{\sf i}\beta$ for $n=0$ which results in singularity at $\beta=1$. For $\beta\to 2$ one obtains the pinch of the contour by two poles at $\nu=-{\sf i}$ and at $\nu={\sf i}-{\sf i}\beta$ for $n=1$ and so on.
@@
For integer $\lambda$ the function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};\lambda)=(R_{d})^{\lambda}\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};0)$ is obtained from the function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};0)$ by applying the operator $(R_{d})^{\lambda}$, which acts on $z$ and $\bar{z}$ variables. It is easy to see that operator $(R_{d})^{\lambda}$ annihilates all functions $(z^{n}+\bar{z}^{n})(z\bar{z})^{{\sf i}\nu-n/2}$ for $n=1,2,\ldots,\lambda-1$ so that for the function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};\lambda)$ the corresponding terms in the sum are absent and the pinch of integration contour appears for $\beta=\lambda+1+k$ where $k=0,1,2\ldots$. Roughly speaking, all singularities at the points $\beta=1,2\ldots,\lambda$, which are present in the function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};0)$ are annihilated by the operator $R_{d}^{\lambda}$ so that the function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};\lambda)$ is regular at these points according to the regularity condition $0<\beta<\frac{D}{2}=\lambda+1$.
@@
The expressions for the ladder integrals in dimensions $D=6,8,\ldots$ can be obtained by the application of the operator $R_{d}$ to the corresponding expression for $D=4$ so that the case $D=4$ plays the crucial role. We have the following expression for the ladder integral in $D=4$ and arbitrary integer $\beta$
@@
In what follows we present how this formula can be used in the case of $\beta=1,2$ and then give some remarks about general case $\beta\in\mathbb{Z}_{>2}$.
Case $\beta=1$: Let us start from the simplest example $\beta=1$ in order to show how everything works and to perform some cross-checks by reproducing the results of subsection 3.1. We have in the case $\beta=1$  [^8] To be rigorous, we should note that the order of operators $R_{d}$ and $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}$ should be the opposite. For simplicity we ignore these subtleties and hope that this will not cause any misunderstandings.
@@
At first sight operation $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}\varepsilon^{L+1}$ should produce vanishing result but function $F_{L+1}(0\,,\varepsilon|z)$ contains singular contribution
@@
so the whole expression can be transformed to the form
@@
At the next step we have to apply operator $R_{d}$ and after that put $\varepsilon\to 0$.
It should be noted that this expression essentially coincides with the expression for the ladder diagram in $D=2$ and therefore has a singularity at $\beta=1$. This singularity manifests itself in the previous formula – it is the contribution $\frac{(-1)^{L+1}}{\varepsilon^{L+1}}$ which is singular for $\varepsilon\to 0$. The convergence condition $0<\beta<D/2=1$ is violated in this case so that the origin of the singularity is clear. The ladder diagram in $D=4$ is obtained by application of the operator $R_{d}$ to the expression for the ladder diagram in $D=2$. In $D=4$, the convergence condition $0<\beta<D/2=2$ is satisfied, so the operator $R_{d}$ must annihilate all singularities, and as we can see this is indeed the case. The last term $O(\varepsilon^{L+1})$ is annihilated by the operator $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}$ so that one obtains
@@
Case $\beta=2$: Now we are going to the next example $\beta=2$. The convergence condition $0<\beta<D/2$ is valid starting from $D=6$ so that we put $\lambda=2$ and get from (3.34)
@@
where in the last line we have removed two contributions. First of all, it is the singular contribution of the form $\frac{1}{\varepsilon^{L+1}}\varphi(z)\varphi(\bar{z})$ which does not contribute to the final answer because it is annihilated by the operator $R_{d}^{2}$. The second contribution is $O(\varepsilon^{L+1})$ and it is annihilated by the operator $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}$. Collecting all factors together we obtain the following expression
@@
Case $\beta\in\mathbb{Z}_{>2}$: The generalization of the above calculations is rather straightforward. For $\beta\in\mathbb{Z}_{>2}$ the convergence condition $0<\beta<D/2$ is fulfilled starting from $D=2\beta+2$ so that $\lambda=\beta$ and one obtains the following representation for the (3.34)
@@
where we omitted all the contributions annihilated by the operator $R_{d}^{\beta}$ and $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}$, and polynomial $\varphi_{\beta}(z)$ is defined by the formula
@@ Conclusions
In this work, we have studied the family of conformal four-point ladder diagrams. This series of integrals was first calculated in the four-dimensional case with physical propagator powers [60, 61] and has continued to attract attention due to the rich underlying structure. The generalized class of four-dimensional diagrams was later shown to have remarkable analytic properties [62]. The generalization to the higher even dimensions turned out to be suitable for application of the theory of graphical functions [86], with the explicit result obtained in [88]. Beyond their interest for computational challenges, these integrals were found to be connected with twisted partition functions [73, 74, 75]. Significant progress was also made in understanding the recursive relations connecting diagrams in different dimensions and with different loop numbers [76].
@@
Using the iterative structure of the ladder diagrams, we showed in [70] that the graph-building operator method together with conformal quantum mechanics provides an explicit answer for the integrals in arbitrary dimension. The present work can be seen as a continuation of those studies. In the paper, we have explicitly derived that the representation obtained in [70] satisfies dimensional and loop shift identities for arbitrary dimensions and general propagator powers described by the parameter $\beta$. Additionally, in two-dimensions and generic $\beta$ we observed a remarkable factorization, analogous to the two-dimensional fishnet diagrams [27], which via dimensional shift identity can be translated to all even dimensions. Specializing further, we showed that for $\beta=1$ (which in the four-dimensional case corresponds to the physical propagator powers) and even $D$ our representation can be rewritten in terms of classical polylogarithms with the rational functions. For higher integer values of $\beta$ the representation based on two-dimensional factorization turned out to be more suitable and was studied. Notably, it naturally provides a regularization for the infrared singularities that arise when $\beta=\frac{D}{2}+k$, with $k\in\mathbb{Z}_{\geq 0}$.
@@
Our results establish clear links between the operator-based construction of [70] and alternative approaches to conformal four-point integrals. This not only enriches the graph-building operator method but also offers potential benefits for alternative methods, particularly given the well-studied properties of operator $H_{\alpha}$, reflecting the underlying conformal and integrable structures. We hope that these insights may contribute to revealing further internal symmetries, such as e.g. antipodal self-duality [77], and to the study of more general families of conformal integrals [80].

### v2

@@ 4. The case $D=10$ ($\lambda=4$) and indices on the lines $\alpha=\beta=1$, $\gamma=4$:
3.2 Two-dimensional factorization for generic $\beta$
In subsection 2.2 we concluded that the $D$-dimensional conformal ladder diagrams with the generic index $\beta$ for $D>2$ and $D=2$ are respectively presented in the form (2.32) and (2.33). We also explicitly showed that the representation (2.32) admits a dimensional shift with the help of the operator (2.35Loebbert and Stawinski (2024); Petkou (2021); Karydas et al. (2024, 2025), namely
@@ 3.2 Two-dimensional factorization for generic $\beta$
An amazing consequence of these results is that we can, following the logic used in Loebbert and Stawinski (2024); Karydas et al. (2024); Petkou (2021); Karydas et al. (2025) (and in subsection 2.3 to construct operator (2.47)), reduce the consideration for the arbitrary even dimension case to the two-dimensional case ($\lambda=0$). This can be seen as a huge advantage, since the two-dimensional integrals are known to have a very constraint form (see e.g. Duhr and Porkert (2024)). In particular, we focus on the results of the work Derkachov et al. (2019), where the notable factorization in $z$ and $\bar{z}$ was shown******Note a similar factorization for the two-dimensional two-loop master diagram Derkachev et al. (2022)..
@@
In the case $\lambda=0$ and generic $\beta$, the following representations for the function $\widetilde{\Phi}^{(\beta)}_{L}(z,\bar{z};0)$ hold††††††In the work Loebbert and Stawinski (2024), representation (3.31) is called the Fourier-Mellin representation.:
@@
Proof. The first equality in (3.31) follows from the representation (2.34). To prove (3.32), we use the result from Appendix D
@@
which leads to  (3.32). ∎
Remark 10. Combining the result of proposition 1 (see eq. (2.36)) and representation (3.32), we obtain the result for arbitrary $\lambda\in\mathbb{Z}_{>0}$ and generic $\beta$ based on the two-dimensional factorized formula
@@
Remark 11. In the case $\beta\in\mathbb{Z}$, the factor contained sines can be simplified to exclude the dependence on $\varepsilon$, namely,
@@
3.3 Consideration of conformal ladder diagrams for $\beta=2,3,\ldots$
@@
In this subsection we consider the case of special discrete points $\beta=2,3,\ldots$. The expression $\Phi^{(\beta)}_{L}({\sf x},{\sf y})$ (2.16) for the ladder diagram is a function of the complex variable $\beta$ and this function is defined in a standard way as an analytical continuation from the region $0<\operatorname{Re}\beta<D/2$ where all integrals in $\Phi^{(\beta)}_{L}({\sf x},{\sf y})$ converge absolutely. This region is determined by two conditions. The horizontal line on Fig. 2 has index $D/2-\beta$ and the corresponding singularity of the integrand is integrable under the condition $D/2-\operatorname{Re}\beta<D/2$, which results in $\operatorname{Re}\beta>0$. The vertical line has index $\beta$ and the singularity is integrable under the condition $\operatorname{Re}\beta<D/2$. Combining these conditions we obtain the region of convergence $0<\operatorname{Re}\beta<D/2$. In this region of complex variable $\beta$ the function $\Phi^{(\beta)}_{L}({\sf x},{\sf y})$ is regular. The analytical continuation from this region is a meromorphic function that has poles at the points $\beta=D/2+k=\lambda+1+k$, where $k=0,1,2,\ldots$.
@@
Let us start with the case $D=2$ and use (2.33) to demonstrate the origin of singularities at a discrete set of points $\beta=1,2,3,\ldots$. The contour of integration (see Remark 2) separates two series of poles of the integrand: the first sequence at the points $\nu_{k}=-{\sf i}\frac{n+1}{2}-{\sf i}k,k=0,1,\ldots$ and the second sequence at the points $\nu_{k}=-{\sf i}\beta+{\sf i}\frac{n+1}{2}+{\sf i}k,k=0,1,\ldots$. For real $\beta$ and $0<\beta<1$ such a contour exists for all $n\in\mathbb{Z}_{\geq 0}$. But for $\beta\to 1$ one obtains the pinch of the contour by two poles at $\nu=-{\sf i}\frac{1}{2}$ and at $\nu={\sf i}\frac{1}{2}-{\sf i}\beta$ for $n=0$ which results in singularity at $\beta=1$. For $\beta\to 2$ one obtains the pinch of the contour by two poles at $\nu=-{\sf i}$ and at $\nu={\sf i}-{\sf i}\beta$ for $n=1$ and so on.
@@
For integer $\lambda$ the function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};\lambda)=(R_{d})^{\lambda}\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};0)$ is obtained from the function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};0)$ by applying the operator $(R_{d})^{\lambda}$, which acts on the $z$ and $\bar{z}$ variables. It is easy to see that the operator $(R_{d})^{\lambda}$ annihilates all functions $(z^{n}+\bar{z}^{n})(z\bar{z})^{{\sf i}\nu-n/2}$ for $n=1,2,\ldots,\lambda-1$ so that for the function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};\lambda)$ the corresponding terms in the sum are absent and the pinch of integration contour appears for $\beta=\lambda+1+k$, where $k=0,1,2\ldots$. Roughly speaking, all singularities at the points $\beta=1,2\ldots,\lambda$, which are present in the function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};0)$, are annihilated by the operator $R_{d}^{\lambda}$ so that the function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};\lambda)$ is regular at these points according to the regularity condition $0<\beta<\frac{D}{2}=\lambda+1$.
@@
The expressions for the ladder integrals in dimensions $D=6,8,\ldots$ can be obtained by the application of the operator $R_{d}$ to the corresponding expression for $D=4$ so that the case $D=4$ plays the crucial role. We have the following expression for the ladder integral in $D=4$ and arbitrary integer $\beta$:
@@
In what follows we present how this formula can be used in the case of $\beta=1,2$ and then give some remarks about the general case $\beta\in\mathbb{Z}_{>2}$.
Case $\beta=1$: Let us start from the simplest example $\beta=1$ in order to show how everything works and to perform some cross-checks by reproducing the results of subsection 3.1. We have in the case $\beta=1$ ‡‡‡‡‡‡To be rigorous, we should note that the order of the operators $R_{d}$ and $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}$ should be opposite. For simplicity we ignore these subtleties and hope that this will not cause any misunderstanding.
@@
At first sight, the operation $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}\varepsilon^{L+1}$ should produce a vanishing result, but the function $F_{L+1}(0\,,\varepsilon|z)$ contains a singular contribution
@@
so the whole expression can be transformed to
@@
At the next step, we have to apply the operator $R_{d}$ and after that put $\varepsilon\to 0$.
It should be noted that this expression essentially coincides with the expression for the ladder diagram in $D=2$ and, therefore, has a singularity at $\beta=1$. This singularity manifests itself in the previous formula – it is the contribution $\frac{(-1)^{L+1}}{\varepsilon^{L+1}}$ which is singular as $\varepsilon\to 0$. The convergence condition $0<\beta<D/2=1$ is violated in this case so that the origin of the singularity is clear. The ladder diagram in $D=4$ is obtained by applying the operator $R_{d}$ to the expression for the ladder diagram in $D=2$. In $D=4$, the convergence condition $0<\beta<D/2=2$ is satisfied, so the operator $R_{d}$ must annihilate all singularities, and as we can see this is indeed the case. The last term $O(\varepsilon^{L+1})$ is annihilated by the operator $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}$ so that one obtains
@@
Case $\beta=2$: Now we proceed to the next example $\beta=2$. The convergence condition $0<\beta<D/2$ is valid starting from $D=6$ so that we put $\lambda=2$ and get from (3.34)
@@
where in the last line we have removed two contributions. First of all, it is the singular contribution of the form $\frac{1}{\varepsilon^{L+1}}\varphi(z)\varphi(\bar{z})$ which does not contribute to the final answer because it is annihilated by the operator $R_{d}^{2}$. The second contribution is $O(\varepsilon^{L+1})$ and it is annihilated by the operator $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}$. Collecting all factors together we obtain the following expression:
@@
Case $\beta\in\mathbb{Z}_{>2}$: The generalization of the above calculations is rather straightforward. For $\beta\in\mathbb{Z}_{>2}$ the convergence condition $0<\beta<D/2$ is fulfilled starting from $D=2\beta+2$ so that $\lambda=\beta$ and one obtains the following representation for the (3.34):
@@
where we omitted all the contributions annihilated by the operator $R_{d}^{\beta}$ and $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}$, and the polynomial $\varphi_{\beta}(z)$ is defined by the formula
@@ 4 Conclusions
In this work, we have studied the family of conformal four-point ladder diagrams. This series of integrals was first calculated in the four-dimensional case with physical propagator powers Usyukina and Davydychev (1993a, b) and continued to attract attention due to the rich underlying structure. The generalized class of four-dimensional diagrams was later shown to have remarkable analytic properties Drummond (2013). The generalization to higher even dimensions turned out to be suitable for application of the theory of graphical functions Borinsky and Schnetz (2022a) with the explicit result obtained in Borinsky and Schnetz (2022b). Besides their interest as computational challenges, these integrals were found to be connected with twisted partition functions Petkou (2021); Karydas et al. (2024, 2025). Significant progress was also made in understanding the recursive relations connecting diagrams in different dimensions and with different loop numbers Loebbert and Stawinski (2024).
@@
Using the iterative structure of the ladder diagrams, we showed in Derkachov et al. (2023) that the graph-building operator method together with conformal quantum mechanics provides an explicit answer for the integrals in arbitrary dimension. The present work can be seen as a continuation of those studies. In the paper, we have explicitly derived that the representation obtained in Derkachov et al. (2023) satisfies dimensional and loop shift identities for arbitrary dimensions and general propagator powers described by the parameter $\beta$. Additionally, in two-dimensions and generic $\beta$ we observed a remarkable factorization analogous to the two-dimensional fishnet diagrams Derkachov et al. (2019), which by the dimensional shift identity can be translated to all even dimensions. Specializing further, we showed that for $\beta=1$ (which in the four-dimensional case corresponds to the physical propagator powers) and even $D$ our representation can be rewritten in terms of classical polylogarithms with the rational functions. For higher integer values of $\beta$ the representation based on two-dimensional factorization turned out to be more suitable and was studied. Notably, it naturally provides a regularization for the infrared singularities that arise when $\beta=\frac{D}{2}+k$ with $k\in\mathbb{Z}_{\geq 0}$.
@@
Our results establish clear links between the operator-based construction of Derkachov et al. (2023) and alternative approaches to conformal four-point integrals. This not only enriches the graph-building operator method but also offers potential benefits for alternative methods, particularly given by the well-studied properties of the operator $H_{\alpha}$, reflecting the underlying conformal and integrable structures. We hope that these insights may contribute to revealing further internal symmetries such as, e.g., antipodal self-duality Dixon and Duhr (2025), and studying of more general families of conformal integrals He and Jiang (2025).
Acknowledgments